# Quantum simulation of strong Charge-Parity violation and Peccei-Quinn mechanism

### v1

@@ A Legendre transformation gives the Hamiltonian form
The Peccei-Quinn mechanism. To dynamically relax the $CP$-violating angle, we promote $\theta$ to a dynamical variable by introducing an axion field $a$ with coupling scale $f_{a}$, such that
$\theta\;\rightarrow\;\theta_{\rm eff}=\theta+\frac{a}{f_{a}}.$ (20)
The gauge-axion interaction term then becomes
$H_{\theta+a}=\frac{g^{2}}{2}\!\left(E-\frac{\theta_{\rm eff}}{2\pi}\right)^{\!2}.$ (21)
The scale $f_{a}$ controls how strongly the axion couples to the gauge sector, where large $f_{a}$ corresponds to weak back-action and slow relaxation of $\theta_{\rm eff}$, while smaller $f_{a}$ enhances the axion response.
The axion degree of freedom can represented as a periodic rotor with compact potential
$H_{a}=-\kappa\cos\!\left(\theta+\frac{a}{f_{a}}\right),$ (22)
which explicitly encodes the $2\pi$ periodicity of the axion. Here, $\kappa$ is the amplitude of the periodic axion potential, which determines how tightly the axion is pulled toward the $CP$-conserving minimum.
@@
Finally, the full axion–coupled Hamiltonian is given by
$H_{\text{total}}=H_{m}+H_{\text{hop}}+H_{\theta+a}+H_{a}+H_{G}.$ (23)
This axion dynamics drive the effective $CP$ phase to $\theta_{\mathrm{eff}}=\theta+a/f_{a}$, realizing the dynamical cancellation of $CP$ violation.
@@
Numerical simulation. We compute the ground state of the two-site Schwinger-model Hamiltonian in its qubit representation using the feedback-based FALQON algorithm [PhysRevLett.129.250502]. FALQON updates the control parameters in real time to ensure a strictly monotonic decrease of the energy expectation value, enabling ground-state preparation without any classical optimization loop. We evaluate the resulting states across different parameter regimes to examine confinement and $CP$-violating effects. For benchmarking and calibration, we focus on three representative configurations:
@@
    Near-continuum: $w=1$, $g^{2}=0.5$, $m=0.25$, and $\lambda=40$. In this limit, the system approaches the weak-coupling or near-continuum regime. Increasing the ratio $w/g$ (i.e., decreasing $g$) allows one to track how the ground-state energy and charge distribution evolve with reduced discretization effects.
@@
    Balanced benchmark: $w=1$, $g^{2}=1$, $m=0.25$, and $\lambda=40$. This choice provides a well-balanced regime where fermionic tunneling, gauge energy, and mass terms contribute comparably, yielding a nontrivial but numerically stable energy spectrum.
@@
    Strong gauge coupling: $w=1$, $g^{2}=2$, $m=0.25$, and $\lambda=40$. Increasing $g^{2}$ amplifies the electric-field contribution to the energy, which reduces charge fluctuations and favors configurations with minimal local net charge.
@@
For the Gauss-law constraint, we use $c_{0}=c_{1}=-\tfrac{1}{2},c_{\ell}=0,d_{0\ell}=+\tfrac{1}{2},d_{1\ell}=-\tfrac{1}{2}.$ For the axion sector, we set $\kappa$ and $f_{a}$ to be comparable to the gauge-field energy scale, choosing $\kappa=g^{2},f_{a}=2.$ The full parameter set is summarized in Tab. 1.
Model $w$ $g^{2}$ $m$ $\lambda$ Gauss-law Axion
Near-continuum 1 0.5 0.25 40 $\begin{aligned} c_{0}&=c_{1}=-\tfrac{1}{2},\\[-2.84526pt] c_{\ell}&=0,\\[-2.84526pt] d_{0\ell}&=+\tfrac{1}{2},\\[-2.84526pt] d_{1\ell}&=-\tfrac{1}{2}\end{aligned}$ $\begin{aligned} \kappa&=g^{2},\\[-2.84526pt] f_{a}&=2\end{aligned}$
Balanced benchmark 1 1 0.25 40
Strong gauge coupling 1 2 0.25 40
Figure 1 shows the ground-state (vacuum) energy $E_{0}(\theta)$ as a function of $\theta$ for three representative parameter regimes: the near-continuum limit (blue circles), the balanced benchmark (orange triangles), and the strong gauge coupling (green triangles).
Figure 1(a) presents the ground-state energy landscape for the system without an axion field, i.e., Eq. (19). In all examined cases, $E_{0}(\theta)$ reaches its minimum at a finite $\theta=2\pi$. This behavior is consistent with theoretical expectations from QCD, where the vacuum energy is generally shifted away from $\theta=0$, indicating a violation of $CP$ symmetry.
Figure 1(b) shows the corresponding results after coupling the system to a dynamical axion mode, described by Eq. (23). The axion modifies the electric-field bias through the shift $\theta\rightarrow\theta_{\mathrm{eff}}=\theta+a/f_{a}$ and introduces an additional potential $H_{a}$ that reshapes the vacuum structure. The resulting ground-state energy retains a $2\pi$-periodic profile, and the global minimum is now located at $\theta=0$ for all examined cases. This behavior reflects the essential Peccei-Quinn mechanism: axion back-reaction dynamically removes the $CP$-violating contribution to the vacuum and drives the system toward $\theta_{\mathrm{eff}}=0$. The restoration of the $CP$-symmetric vacuum also explains why experiments observe the physical value of $\theta$ to lie extremely close to zero [PhysRevLett.97.131801, PhysRevLett.124.081803].
These results show that even a minimal, few-qubit realization of the axion-Schwinger model captures the essential features of topological gauge interactions. As seen from the contrast between Fig. 1(a) and Fig. 1(b), the model without an axion settles at a nonzero value of $\theta$, consistent with the theoretical expectation from QCD that a generic $\theta$-term leads to $CP$ violation. When the axion field is introduced, the system instead relaxes to $\theta=0$, illustrating how the dynamical axion mechanism naturally resolves the strong $CP$ problem.
Figure 1: Vacuum energy of the two sites lattice Schwinger model with and without a dynamical axion. (a) Ground (vacuum) energy $E_{0}(\theta)$ of the qubit-encoded Schwinger model for three representative parameter regimes. In all cases, the minimum appears at a finite, nonzero value of $\theta$, consistent with the theoretical expectation that a generic $\theta$-term leads to CPCP violation in strong interactions. (b) Ground (vacuum) energy after coupling the model to a dynamical axion field. The axion back-reaction reshapes the $\theta$-dependent potential, producing a $2\pi$-periodic structure whose global minimum is pinned at $\theta=0$. This shift from a displaced vacuum in (a) to a $CP$-symmetric vacuum in (b) shows that the axion-Schwinger construction reproduces the dynamical relaxation mechanism central to the Peccei-Quinn solution of the strong $CP$ problem.
@@
Extended Lattice: 4 Sites and 3 Links (7 Qubits). Figure 2 summarizes the ground-state energy landscape of the qubit-encoded Schwinger model extended to a four-site lattice with three gauge links. The explicit Hamiltonian, including mass, gauge-mediated hopping, electric-field energy, and Gauss-law terms, is given in the End Matter section. As in Fig. 1, we evaluate the vacuum energy $E_{0}(\theta)$ for three representative parameter regimes: near-continuum (blue), balanced benchmark (orange), and strong gauge coupling (green).
Figure 2(a) shows the case without the axion field. Similar as the two-site setup of Fig. 1(a), the enlarged lattice exhibits a displaced vacuum minimum at nonzero $\theta=\pi$. This behavior reflects a finite-volume effect, where in an increased number of link degrees of freedom modifies the electric-field distribution and can stabilize a nonvanishing topological angle rather than a $CP$-symmetric configuration.
@@
Figure 2(b) displays the same model after coupling to a dynamical axion. As in Fig. 1(b), the axion shifts the effective vacuum angle $\theta\rightarrow\theta_{\mathrm{eff}}$ and introduces a periodic restoring potential. The resulting energy profile becomes explicitly $2\pi$-periodic, and its global minimum is driven back to $\theta=0$ for all parameter regimes. This dynamical relaxation embodies the Peccei-Quinn mechanism where axion back-reaction cancels the effective $CP$ phase and restores a $CP$-symmetric vacuum, even in a minimal seven-qubit realization.
@@
Figure 2: Vacuum energy of the four sites lattice Schwinger model with and without a dynamical axion. Ground-state (vacuum) energy $E_{0}(\theta)$ of the qubit-encoded Schwinger model for three representative parameter regimes, shown using the same conventions as Fig. 1 for (a) without axion field, and (b) with axion field.
@@
Conclusion. Starting from the QCD Lagrangian with a $CP$-violating $\theta$-term, we derived its Hamiltonian and showed how the $CP$-odd contribution alters the canonical gauge structure. Reducing the theory to its $(1+1)$-dimensional Schwinger analogue yields a minimal model that retains key nonperturbative features, including topological vacuum structure and $\theta$-sensitivity. Using the Jordan-Wigner and quantum-link mappings, we express this model as an explicit Pauli Hamiltonian suitable for digital quantum simulation. We then prepared its ground state via the FALQON algorithm and evaluated the vacuum energy as a function of $\theta$, confirming the displaced vacuum expected from strong interactions. When a dynamical axion field is introduced, the vacuum relaxes to $\theta=0$, providing a direct quantum-simulation realization of the Peccei-Quinn mechanism and a controlled framework to study $CP$ violation and its dynamical cancellation.
@@
I End Matter
Jordan-Wigner Transformation. To simulate the (1+1)(1+1)D lattice Schwinger model on a quantum computer, the Hamiltonian must be expressed entirely in terms of Pauli matrices acting on qubits. This requires mapping both the fermionic matter fields and the gauge link variables to spin-$\tfrac{1}{2}$ operators. The fermionic fields are encoded using the Jordan–Wigner (JW) transformation, while the gauge links are represented by spin operators in the quantum link model (QLM), which provides a finite-dimensional truncation of the continuous gauge degrees of freedom.
Fermionic mapping.
@@ I End Matter
$\psi_{x}$ $=\Big(\prod_{y<x}Z_{y}\Big)S_{x}^{-},$ $\psi_{x}^{\dagger}$ $=\Big(\prod_{y<x}Z_{y}\Big)S_{x}^{+},$ (24)
@@
The electric-field energy, including the $\theta$-dependent $CP$-violating shift, takes the form
$H_{\theta}$ $=\frac{g^{2}}{2}\Big(E-\frac{\theta}{2\pi}\Big)^{2}$
$=\frac{g^{2}}{2}\!\left[\!\left(\frac{\theta}{2\pi}\right)^{2}-\frac{\theta}{2\pi}+\frac{1}{2}\!\right]I_{\ell}+\frac{g^{2}}{2}\!\left(\frac{\theta}{2\pi}-\frac{1}{2}\right)\!Z_{\ell}.$ (28)
@@
which involves correlated three-body interactions between the two matter qubits and the gauge-link qubit, ensuring local gauge invariance.
Gauss’s law enforcement.
Local charge conservation is imposed energetically through a penalty term
@@
$H_{\text{total}}=H_{m}+H_{\text{hop}}+H_{\theta}+H_{G},$ (31)
where $H_{m}$ encodes the fermionic mass asymmetry, $H_{\text{hop}}$ the correlated fermion–gauge tunneling, $H_{\theta}$ the $CP$-violating electric-field shift, and $H_{G}$ enforces gauge invariance. Despite its minimal size, this model captures essential features of the Schwinger mechanism and provides a compact benchmark for near-term quantum simulations of gauge-theory dynamics.
Ground-state preparation via FALQON.
To obtain the ground state of the lattice Schwinger Hamiltonian, we employ the Feedback-based Algorithm for Quantum Optimization (FALQON) [PhysRevLett.129.250502]. FALQON formulates the search for the ground state as a closed-loop quantum control problem. The system evolves under a time-dependent Hamiltonian composed of a cost term $H_{C}$ (the problem Hamiltonian) and a driver term $H_{D}$ that induces state transitions:

### v2

@@ Abstract
The Peccei-Quinn mechanism. To dynamically relax the $CP$-violating angle, we promote the parameter $\bar{\theta}$ to a dynamical variable by introducing an axion field $a(x)$ with decay constant $f_{a}$. At low energies, the effective QCD-axion Lagrangian is
$\mathcal{L}_{\rm QCD+a}=\mathcal{L}_{\rm QCD}\big|_{\bar{\theta}\to\bar{\theta}+\frac{a}{f_{a}}}+\frac{1}{2}\partial_{\mu}a\partial^{\mu}a-V\left(\bar{\theta}+\frac{a}{f_{a}}\right).$ (20)
The axion enters as a shift of the effective $CP$-violating angle, $\theta_{\rm eff}\equiv\bar{\theta}+a/f_{a}$, rendering it dynamical.
At energies below the QCD confinement scale, the axion potential is well described in a model-independent way by the topological susceptibility $\chi$,
$V(\theta_{\rm eff})\simeq\chi[1-\cos(\theta_{\rm eff})].$ (21)
Expanding around the minimum yields the standard axion mass relation $m_{a}^{2}f_{a}^{2}=\chi$.
The axion Hamiltonian density follows from the canonical momentum $\pi_{a}=\dot{a}$ and is given by
$\mathcal{H}_{a}=\frac{1}{2}\pi_{a}^{2}+\frac{1}{2}(\partial_{x}a)^{2}+V(\theta_{\rm eff}),$ (22)
with the total axion Hamiltonian $H_{a}=\int dx\mathcal{H}_{a}$. The coupling between the axion and the gauge sector appears through the shifted $\bar{\theta}$ term,
$H_{\bar{\theta}+a}=\frac{g^{2}}{2}\left(E-\frac{\theta_{\rm eff}}{2\pi}\right)^{2}.$ (23)
@@
Combining all contributions, the full Hamiltonian of the axion-coupled system reads
$H_{\rm total}=H_{m}+H_{\rm hop}+H_{\bar{\theta}+a}+H_{a}+H_{G}.$ (24)
The axion dynamics drive the system toward the minimum of the effective potential. Minimization with respect to $a$ enforces $\langle a\rangle=-f_{a}\bar{\theta}$, so that the effective angle relaxes to $\theta_{\rm eff}=0$. This realizes the Peccei-Quinn mechanism at the Hamiltonian level: the axion back-reaction dynamically cancels the $CP$-violating phase, restoring a $CP$-symmetric vacuum.
@@
Numerical simulation. We compute the ground state of the two-site Schwinger-model Hamiltonian in its qubit representation using the feedback-based FALQON algorithm [17]. FALQON updates the control parameters in real time to ensure a strictly monotonic decrease of the energy expectation value, enabling ground-state preparation without any classical optimization loop. See detail of FALQON in App. B.
We evaluate the resulting states across different parameter regimes to examine confinement and $CP$-violating effects. For benchmarking and calibration, we focus on three representative configurations:
Model $w$ $g^{2}$ $m$ $\lambda$ Gauss-law Axion
Near-continuum 1 1 0.25 40 $\begin{aligned} c_{0}&=-\tfrac{1}{2},\\ c_{1}&=\tfrac{1}{2},\\ c_{\ell}&=0,\\ d_{0\ell}&=d_{1\ell}=-\tfrac{1}{2}\end{aligned}$ $\begin{aligned} m_{a}&=g^{2},\\[-2.84526pt] f_{a}&=2\end{aligned}$
Balanced benchmark 1 2 0.25 40
Strong gauge coupling 1 5 0.25 40
Figure 1: Vacuum energy of the two-site lattice Schwinger model with and without a dynamical axion. (a) Vacuum energy $E_{0}(\bar{\theta})$ as a function of $\bar{\theta}$ for three representative parameter regimes. In the $(1+1)$-dimensional lattice formulation, reduced dimensionality and discretization effects lift intermediate branches, such that only the minima at $\bar{\theta}=0$ and $2\pi$ are clearly resolved. (b) Vacuum energy after coupling to a dynamical axion field. The axion back-reaction drives the effective angle $\theta_{\rm eff}=\bar{\theta}+a/f_{a}$ toward zero, rendering $E_{0}$ independent of $\bar{\theta}$. Inset: Schematic axion-induced effective potential illustrating the relaxation of $\theta_{\rm eff}$ toward its global minimum.
@@
    Near-continuum: $w=1$, $g^{2}=1$, $m=0.25$, and $\lambda=40$. In this limit, the system approaches the weak-coupling or near-continuum regime. Increasing the ratio $w/g$ (i.e., decreasing $g$) allows one to track how the ground-state energy and charge distribution evolve with reduced discretization effects.
@@
    Balanced benchmark: $w=1$, $g^{2}=2$, $m=0.25$, and $\lambda=40$. This choice provides a well-balanced regime where fermionic tunneling, gauge energy, and mass terms contribute comparably, yielding a nontrivial but numerically stable energy spectrum.
@@
    Strong gauge coupling: $w=1$, $g^{2}=5$, $m=0.25$, and $\lambda=40$. Increasing $g^{2}$ amplifies the electric-field contribution to the energy, which reduces charge fluctuations and favors configurations with minimal local net charge.
@@
For the Gauss-law constraint, we use $c_{0}=-\tfrac{1}{2},c_{1}=\tfrac{1}{2},c_{\ell}=0,d_{0\ell}=d_{1\ell}=-\tfrac{1}{2}.$ For the axion sector, we set $m_{a}$ and $f_{a}$ to be comparable to the gauge-field energy scale, choosing $m_{a}=g^{2},f_{a}=2.$ The full parameter set is summarized in Tab. 1.
Fig. 1 (a) shows the vacuum energy $E_{0}(\bar{\theta})$ as a function $\bar{\theta}$ (in units of $\pi)$ for three representative parameter regimes: near-continuum, balanced benchmark, and strong gauge coupling, in the absence of an axion field [Hamiltonian Eq. (19)]. In the continuum QCD, $E_{0}(\bar{\theta})$ is expected to be a $2\pi$-periodic function of $\bar{\theta}$, with degenerate minima at $\bar{\theta}=0\ (\mathrm{mod}\ 2\pi)$. In the present (1+1)(1+1)dimensional Schwinger-model approximation, lattice discretization, reduced dimensionality, and strong-coupling effects distort this periodic structure. Consequently, only the minima at $\bar{\theta}=0$ and $2\pi$ are clearly resolved, while intermediate branches are lifted. This distortion becomes more pronounced at stronger gauge coupling, where the vacuum energy exhibits a steeper curvature away from the minima.
Fig. 1 (b) illustrates the effect of coupling the system to a dynamical axion field. Starting from a generic initial value of $\bar{\theta}$, the axion dynamically relaxes toward the minimum of the combined potential, driving the effective angle $\theta_{\rm eff}=\bar{\theta}+a/f_{a}$ toward zero. As a result, the relaxed vacuum energies become independent of $\bar{\theta}$, demonstrating the dynamical cancellation of $CP$ violation. The inset schematically illustrates the axion-induced effective potential as a function of $\bar{\theta}$, where the axion field shifts $\theta_{\rm eff}$ toward the global minimum. Despite dimensional and lattice artifacts inherent to the (1+1)(1+1)D formulation, these results confirm that the Peccei-Quinn mechanism remains operative at the Hamiltonian level.
@@
Extended lattice: 4 sites and 3 links (7 qubits). Figure 2 summarizes the ground-state energy landscape of the qubit-encoded Schwinger model extended to a four-site lattice with three gauge links. The explicit Hamiltonian, including mass, gauge-mediated hopping, electric-field energy, and Gauss-law terms, is given in App. C. As in Fig. 1, we evaluate the vacuum energy $E_{0}(\bar{\theta})$ for three representative parameter regimes: near-continuum (blue), balanced benchmark (orange), and strong gauge coupling (green).
Figure 2(a) shows the case without the axion field. Similar to the two-site case shown in Fig. 1(a), the vacuum energy exhibits minima at $\bar{\theta}=0$ and $2\pi$. In the enlarged lattice, the increased number of gauge-link degrees of freedom modifies the electric-field distribution, leading to finite-volume effects that distort the ideal continuum periodic structure. As a result, the vacuum does not exhibit a fully resolved $2\pi$-periodic landscape, and intermediate branches are lifted relative to the minima.
@@
Figure 2(b) displays the same model after coupling to a dynamical axion. As in Fig. 1(b), the axion dynamically shifts the effective angle $\bar{\theta}\rightarrow\theta_{\rm eff}=\bar{\theta}+a/f_{a}$ and introduces a restoring potential. After relaxation, the vacuum energy becomes independent of $\bar{\theta}$, with the global minimum driven to $\theta_{\rm eff}=0$ for all parameter regimes. This behavior demonstrates that the Peccei-Quinn mechanism remains operative and robust against finite-volume effects, restoring a $CP$-symmetric vacuum even in an enlarged lattice realization implemented with seven qubits.
@@
Figure 2: Vacuum energy of the four sites lattice Schwinger model with and without a dynamical axion. Vvacuum energy $E_{0}(\bar{\theta})$ of the qubit-encoded Schwinger model for three representative parameter regimes, shown using the same conventions as Fig. 1 for (a) without axion field, and (b) with axion field.
@@
Conclusion. Starting from the QCD Lagrangian with a $CP$-violating $\bar{\theta}$ term, we derived the corresponding Hamiltonian and identified how this term modifies the gauge-sector structure. By reducing the theory to its $(1+1)$-dimensional Schwinger-model analogue, we obtained a minimal setting that preserves essential nonperturbative features, including topological vacuum structure and sensitivity to $\bar{\theta}$. Using Jordan-Wigner and quantum-link mappings, the model was rewritten as an explicit Pauli Hamiltonian suitable for digital quantum simulation. We evaluated the ground state using the FALQON algorithm and computed the vacuum energy as a function of $\bar{\theta}$, reproducing the expected vacuum behavior of a theory with a $CP$-violating term. Upon coupling the system to a dynamical axion field, the vacuum energy relaxes to a minimum at $\theta_{\rm eff}=0$, demonstrating the dynamical cancellation of $CP$ violation. These results provide a concrete quantum-simulation realization of the Peccei-Quinn mechanism and establish a controlled platform for studying $CP$ violation and its dynamical relaxation in gauge theories.
@@
Appendix A Appendix A: Jordan-Wigner Transformation
To simulate the (1+1)(1+1)D lattice Schwinger model on a quantum computer, the Hamiltonian must be expressed entirely in terms of Pauli matrices acting on qubits. This requires mapping both the fermionic matter fields and the gauge link variables to spin-$\tfrac{1}{2}$ operators. The fermionic fields are encoded using the Jordan-Wigner (JW) transformation, while the gauge links are represented by spin operators in the quantum link model (QLM), which provides a finite-dimensional truncation of the continuous gauge degrees of freedom.
@@ Appendix A Appendix A: Jordan-Wigner Transformation
$\psi_{x}=\Big(\prod_{y<x}Z_{y}\Big)S_{x}^{-},\qquad\psi_{x}^{\dagger}=\Big(\prod_{y<x}Z_{y}\Big)S_{x}^{+},$ (25)
@@
The electric-field energy, including the $\bar{\theta}$-dependent $CP$-violating shift, takes the form
$H_{\bar{\theta}}$ $=\frac{g^{2}}{2}\Big(E-\frac{\bar{\theta}}{2\pi}\Big)^{2}$
$=\frac{g^{2}}{2}\!\left[\!\left(\frac{\bar{\theta}}{2\pi}\right)^{2}-\frac{\bar{\theta}}{2\pi}+\frac{1}{2}\!\right]I_{\ell}+\frac{g^{2}}{2}\!\left(\frac{\bar{\theta}}{2\pi}-\frac{1}{2}\right)\!Z_{\ell}.$ (29)
@@
which involves correlated three-body interactions between the two matter qubits and the gauge-link qubit, ensuring local gauge invariance. Local charge conservation is imposed energetically through a penalty term
@@
$H_{\text{total}}=H_{m}+H_{\text{hop}}+H_{\bar{\theta}}+H_{G},$ (32)
where $H_{m}$ encodes the fermionic mass asymmetry, $H_{\text{hop}}$ the correlated fermion-gauge tunneling, $H_{\bar{\theta}}$ the $CP$-violating electric-field shift, and $H_{G}$ enforces gauge invariance. Despite its minimal size, this model captures essential features of the Schwinger mechanism and provides a compact benchmark for near-term quantum simulations of gauge-theory dynamics.
Appendix B Appendix B: Ground-state preparation via FALQON
To obtain the ground state of the lattice Schwinger Hamiltonian, we employ the Feedback-based Algorithm for Quantum Optimization (FALQON) [17]. FALQON formulates the search for the ground state as a closed-loop quantum control problem. The system evolves under a time-dependent Hamiltonian composed of a cost term $H_{C}$ (the problem Hamiltonian) and a driver term $H_{D}$ that induces state transitions: